Calculate the density of nitrogen gas at 227c and 5atm pressure if its density at STP is 0.4g/ml
 
 

Grade:12th pass

1 Answers

Arun
5 years ago
It is significant to know that the P is a pressure, Vis a volume, m is the moles and R is a universal gas constant.
 The value of R is = 0.0821 liter*atm/ (mol*K).  
When it comes to T, it is a thermodynamic temperature.  
T = 227 + 273 =500 k 
D = pm ÷ RT
D = 5 × 28 ÷ 0.0821 × 500
D = 3.41

Regards
Arun